HETAV PATEL

Grade 11,

Find the common factors of the given terms. 2y, 22xy

Find the common factors of the given terms. 
2y, 22xy

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Factors of 2y and 22xy
2y = 2 × y
22xy = 2 × 11 × x × y
Common factors of 2y and 22xy are 2, y
and ,2×y = 2y

Thanks